# AI-Accelerated Flow Simulation: A Robust Auto-Regressive Framework for Long-Term CFD Forecasting

### v3

@@ Abstract
This study addresses the critical challenge of error accumulation in spatio-temporal auto-regressive (AR) predictions within scientific machine learning models by exploring temporal integration schemes and adaptive multi-step rollout strategies. We introduce the first implementation of the two-step Adams-Bashforth method specifically tailored for data-driven AR prediction, leveraging historical derivative information to enhance numerical stability without additional computational overhead. To validate our approach, we systematically evaluate time integration schemes across canonical 2D PDEs (advection, heat, and Burgers’ equations) before extending to complex Navier-Stokes cylinder vortex shedding dynamics. Additionally, we develop three novel adaptive weighting strategies that dynamically adjust the importance of different future time steps during multi-step rollout training. Our analysis reveals that as physical complexity increases, such sophisticated rollout techniques become essential, with the Adams-Bashforth scheme demonstrating consistent robustness across investigated systems and our best adaptive approach delivering an 89% improvement over conventional fixed-weight methods while maintaining similar computational costs. For the complex Navier-Stokes vortex shedding problem, despite using an extremely lightweight graph neural network with just 1,177 trainable parameters and training on only 50 snapshots, our framework accurately predicts 350 future time steps—a 7:1 prediction-to-training ratio—reducing mean squared error from 0.125 (single-step direct prediction) to 0.002 (Adams-Bashforth with proposed multi-step rollout). Our integrated methodology demonstrates an 83% improvement over standard noise injection techniques and maintains robustness under severe spatial constraints; specifically, when trained on only a partial spatial domain, it still achieves 58% and 27% improvements over direct prediction and forward Euler methods, respectively. Our framework’s model-agnostic design, operating at the fundamental level of AR prediction mechanics, enables direct integration with any neural network architecture without requiring model-specific modifications, introducing a versatile solution for robust long-term spatio-temporal predictions across various engineering disciplines.
@@ 1 Introduction
2 Methods: time integration schemes and adaptive multi-step rollout
@@ 2.1 Time integration schemes for auto-regressive prediction
A critical aspect of implementing the Adams-Bashforth method in an AR framework is the choice of how to approximate the derivatives $\frac{\delta\mathbf{u}}{\delta t}\big{|}_{t}$ and $\frac{\delta\mathbf{u}}{\delta t}\big{|}_{t-\Delta t}$. We implement the forward Euler method for time integration (Eq. 6 and Eq. 7), a common approach in AI-based temporal prediction [2, 19, 26]. Since both methods employ the forward Euler scheme to compute the derivatives, the forward Euler and Adams-Bashforth approaches yield identical model outputs: see Eq. 1 and Eq. 6. For brevity, we refer to the combined Adams-Bashforth with forward Euler method as “Adams-Euler” throughout this paper.
@@ 2.2 Adaptive multi-step rollout
Multi-step rollout [30] addresses this limitation by incorporating multiple future time steps into the training process: Fig. 2. Specifically, given input snapshots $\mathbf{u}(t),\ldots,\mathbf{u}(t-(N-1))$, the AR framework first predicts $\hat{\mathbf{u}}(t+1)$. The input history is then updated to $\hat{\mathbf{u}}(t+1),\ldots,\mathbf{u}(t-(N-2))$, allowing prediction of $\hat{\mathbf{u}}(t+2)$. This iterative process continues for $M$ steps, producing predictions from $\hat{\mathbf{u}}(t+1)$ to $\hat{\mathbf{u}}(t+M)$ based on the initial input history. The total loss is defined as a weighted sum of individual step losses:
@@
Recognizing that errors often accumulate over time in AR predictions, we propose a simplified approach that focuses only on the first and last steps of the rollout. The adaptive weights are computed as:
@@
By focusing on the first and last steps, this method emphasizes improving both immediate next prediction and long-term prediction, which are critical for model performance. This simplified approach reduces computational complexity and avoids potential overfitting to intermediate steps.
@@ 3.2 Model training details
For each of the 50 samples per PDE type, a separate standard multi-layer perceptron (MLP) model is trained to act as a universal time-stepping function that learns the local temporal dynamics of the system. This model is designed to be independent of spatial coordinates. Specifically, its input is a time-series vector representing the last 30 snapshots of the solution $u$ at a single query point in the domain, and its output is the prediction for that same point’s next state. During both training and inference, the trained MLP is applied simultaneously and independently to every point in the spatial grid, effectively processing all spatial grid points in a large batch. This approach allows the model to learn a general rule for temporal evolution that can be utilized for inference at any spatial location. The final performance metrics are averaged across the 50 independent runs from 50 cases to ensure statistical robustness. The training hyperparameters are kept consistent across all experiments to facilitate a fair comparison:
@@ 3.3 Predictive performance of time integration schemes
We first evaluate the different time integration methods using a single-step rollout ($M=1$). As shown in Table 1, all schemes that predict temporal derivatives consistently outperform the conventional direct prediction method.
@@
PDE Time Integration Scheme Test MSE [×$\times 10^{-6}$] Training Time [s]
@@ 3.4 Application of proposed adaptive multi-step rollout
Single-step Rollout Multi-Step Rollout ($M$=4)
@@ 4.1.1 Dataset and preprocessing
The dataset features a mesh scenario with a cylinder placed in a fluid flow, generating vortex shedding characterized by oscillatory flow patterns. The training mesh consists of 1,946 nodes, 11,208 edges, and 3,658 volume cells, as shown in Fig. 4: this training dataset is adapted from the work of Google DeepMind [19]. The flow conditions include a maximum inlet velocity of $1.78$ m/s with a parabolic velocity distribution and a cylinder diameter of $0.074$ m. A single vortex shedding period comprises approximately 29 snapshots (or trajectories).
For training, we utilize only 50 consecutive $x$-velocity snapshots. Specifically, the model input consists of a sequence of $N=20$ past $x$-velocity snapshots, fewer than the number of snapshots in a vortex shedding period, forming a 20-channel input graph. The task is to predict the flow field of the next immediate snapshot, enabling the model to learn temporal dependencies within the limited training data.
@@ 4.1.2 Model training details
The Graph U-Net architecture employed in this study is designed with a focus on computational efficiency and simplicity, comprising only 1,177 trainable parameters (Fig. 5). The encoder part of the model consists of four graph convolutional network (GCN) layers [33], with channel dimensions decreasing from 20 to 1 in the sequence 20, 15, 10, 5, and 1, where the initial dimension corresponds to the 20-channel input graph. A pooling ratio of 0.6 is applied using the gPool layer [32] after each GCN layer in the encoder, as this ratio was found to be optimal in our previous work [27]. This pooling operation reduces the number of nodes by 60% at each step, effectively capturing multi-scale features while maintaining computational tractability.
@@
In summary, the Graph U-Net model used in this section is evaluated under intentionally harsh conditions, including limited model capacity (1,177 parameters), minimal training data (only 50 snapshots), and a long prediction horizon of 350 future snapshots. These constraints simulate practical engineering scenarios where only small datasets are available and computational infrastructure for model training is limited. This challenging setup rigorously tests the model’s robustness and its ability to manage error accumulation over time, particularly for complex temporal patterns like vortex shedding.
@@ 4.2.1 Predictive performance comparison
Table 3: Comparison of MSE for different time integration schemes over 350 future snapshots. Training time based on NVIDIA 3090 GPU is also shown.
@@
Among the various approaches, Fig. 6 illustrates the fundamental instability of the second-order central difference scheme. Four consecutive snapshots after 100 rollouts reveal highly oscillatory behavior: snapshots after 100 (Fig. 6(a)) and 102 (Fig. 6(c)) rollouts appear similar, while those after 101 (Fig. 6(b)) and 103 (Fig. 6(d)) exhibit similar patterns. This alternating pattern persists beyond 100 rollouts, reflecting the scheme’s fundamental limitation where $\mathbf{u}(t+\Delta t)$ depends heavily on $\mathbf{u}(t-\Delta t)$ while bypassing the intermediate state $\mathbf{u}(t)$; since $\mathbf{u}(t+\Delta t)=\mathbf{u}(t-\Delta t)+2\Delta t\cdot\frac{\delta\mathbf{u}}{\delta t}$. These results align with established findings that central difference schemes are generally unsuitable for time integration due to inherent instability.
@@
To further illustrate performance limitations, Fig. 7 presents flow field predictions after 100 rollout steps for the three outperforming schemes alongside ground truth. The ground truth shows a specific vortex shedding phase behind the cylinder. All methods show noticeable discrepancies from ground truth after only 100 rollout steps. Direct prediction provides time-averaged results where flow fields remain constant during rollout progression. The corresponding error plots highlight significant prediction deviations, especially in the wake region where complex flow dynamics are most pronounced.
@@ 4.3 Extension of time integration schemes into conventional multi-step rollout scenario
We extend the previously investigated time integration schemes by incorporating conventional multi-step rollout techniques with different values of $M$, where $M$ represents the number of future snapshots considered during training. Specifically, we examine $M=1$, $2$, $4$, and $8$. When $M=1$, multi-step rollout is not applied, corresponding to the vanilla AR model. We adopt the conventional weighting strategy suggested by Wu et al. [30], where loss weights are set to $w_{1}=1$ for the first future step and $w_{i}=0.1$ for subsequent steps (see Eq. 8).
@@
Table 4: MSE comparison for different time integration schemes with varying multi-step rollout lengths $M$. Erroneous values are shown with a gray background.
@@
Averaged time [s] 1914 2086 2633 3131
These results reveal two key insights: First, naively increasing $M$ imposes excessive training constraints on the lightweight model (1,177 parameters), often leading to performance degradation despite the theoretical benefit of learning longer-term dependencies. Second, the time integration scheme proves more critical than the choice of $M$—while most methods fail even with optimal $M$ values, Adams-Euler maintains robust performance across all $M$ values, demonstrating inherent stability for long-term AR prediction.
@@
To illustrate the improved performance achieved with Adams-Euler and $M=4$ multi-step rollout (best case in Table 4), Fig. 8 shows flow field predictions after 200 and 300 rollout steps. At snapshot $t+200$ (Figs. 8(a) and 8(b)), the prediction shows reasonable agreement with ground truth, capturing overall flow patterns and vortex shedding behavior. This represents substantial improvement over Fig. 7, where accurate predictions were unattainable even at 100 rollout steps, clearly demonstrating the effectiveness of combining conventional multi-step rollout with the Adams-Euler scheme. However, at snapshot $t+300$ (Figs. 8(c) and 8(d)), MSE increases from 0.091 to 0.105, showing noticeable divergence from ground truth: flow structure accuracy diminishes with different vortex shedding patterns. While the incorporation of multi-step rollout with Adams-Euler significantly enhances long-term prediction performance, room for improvement remains, particularly for prediction horizons extending beyond $t+200$. This motivates the development of adaptive multi-step rollout strategies explored in subsequent sections.

### v4

# Model-Agnostic AI Framework with Explicit Time Integration for Long-Term Fluid Dynamics Prediction

@@ Abstract
The integration of predictive AI into computational design and engineering workflows is often hindered by the challenge of long-term forecasting performance. This study presents a robust framework for AI-accelerated flow simulation, specifically addressing the critical issue of error accumulation in auto-regressive (AR) surrogate models, which is a key bottleneck for their practical use in the design cycle. We introduce the first implementation of the two-step Adams-Bashforth method specifically tailored for data-driven AR prediction, leveraging historical derivative information to enhance numerical stability without additional computational overhead. To validate our approach, we systematically evaluate time integration schemes across canonical 2D PDEs (advection, heat, and Burgers’ equations) before extending to complex Navier-Stokes cylinder vortex shedding dynamics. Additionally, we develop three novel adaptive weighting strategies that dynamically adjust the importance of different future time steps during multi-step rollout training. Our analysis reveals that as physical complexity increases, such sophisticated rollout techniques become essential, with the Adams-Bashforth scheme demonstrating consistent robustness across investigated systems and our best adaptive approach delivering an 89% improvement over conventional fixed-weight methods while maintaining similar computational costs. For the complex Navier-Stokes vortex shedding problem, despite using an extremely lightweight graph neural network with just 1,177 trainable parameters and training on only 50 snapshots, our framework accurately predicts 350 future time steps—a 7:1 prediction-to-training ratio—reducing mean squared error from 0.125 (single-step direct prediction) to 0.002 (Adams-Bashforth with proposed multi-step rollout). Our integrated methodology demonstrates an 83% improvement over standard noise injection techniques and maintains robustness under severe spatial constraints; specifically, when trained on only a partial spatial domain, it still achieves 58% and 27% improvements over direct prediction and forward Euler methods, respectively. Our framework’s model-agnostic design, operating at the fundamental level of AR prediction mechanics, enables direct integration with any neural network architecture without requiring model-specific modifications, introducing a versatile solution for robust long-term spatio-temporal predictions across various engineering disciplines.
@@ 1 Introduction
2 Methods: Explicit Time Integration Schemes and Adaptive Multi-Step Rollout
@@ 2.1 Time integration schemes for auto-regressive prediction
A critical aspect of implementing the Adams-Bashforth method in an AR framework is the choice of how to approximate the derivatives $\frac{\delta\mathbf{u}}{\delta t}\big|_{t}$ and $\frac{\delta\mathbf{u}}{\delta t}\big|_{t-\Delta t}$. We implement the forward Euler method for time derivatives (Eq. 6 and Eq. 7), a common approach in AI-based temporal prediction [2, 19, 26]: for brevity, we refer to this combined Adams-Bashforth with forward Euler method as “Adams-Euler” throughout this paper. Since forward Euler and Adams-Euler methods employ the forward Euler scheme to compute time derivatives, those approaches yield identical model outputs: see Eq. 1 and Eq. 6. Given this equivalence in training, a single trained model could theoretically be used for both forward Euler and Adams-Euler inference. However, for ease of experimental implementation, we simply trained separate models for each scheme. To ensure this approach provides a fair and robust comparison, each experiment was repeated multiple times using different random seeds, and the averaged results are presented throughout the paper. Consequently, the minor variations in reported training times between the forward Euler and Adams-Euler schemes are a natural result of these separate training runs and normal stochastic fluctuations in GPU computation.
@@ 2.2 Adaptive multi-step rollout
Multi-step rollout [30] addresses this limitation by incorporating multiple future time steps into the training process: Figure 2. Specifically, given input snapshots $\mathbf{u}(t),\ldots,\mathbf{u}(t-(N-1))$, the AR framework first predicts $\hat{\mathbf{u}}(t+1)$. The input history is then updated to $\hat{\mathbf{u}}(t+1),\ldots,\mathbf{u}(t-(N-2))$, allowing prediction of $\hat{\mathbf{u}}(t+2)$. This iterative process continues for $M$ steps, producing predictions from $\hat{\mathbf{u}}(t+1)$ to $\hat{\mathbf{u}}(t+M)$ based on the initial input history. The total loss is defined as a weighted sum of individual step losses:
@@
Recognizing that error accumulation is the primary failure mode in AR prediction, we frame the multi-step learning task as a trajectory control problem. The goal is to enforce both short-term accuracy and long-term stability. To this end, we propose a simplified and robust weighting strategy that focuses only on the two most critical points of the rollout trajectory: the beginning and the end.
  • First step ($w_{1}$): This term enforces short-term accuracy. A precise prediction for the very next time step is crucial, as any initial error will immediately propagate and compound, causing the trajectory to diverge quickly from the ground truth. This penalty ensures the forecast begins on the correct path, translating the momentum of the historical data into the very first predicted step.
  • Last step ($w_{M}$): This term enforces long-term stability. By penalizing the error at the final point of the rollout window, the model is explicitly regularized against divergence. This loss acts as a constraint on the future, forcing the model to learn dynamics that remain stable over the entire horizon.
This approach is motivated by curriculum learning [31], an AI training strategy where a model is first taught simpler concepts before progressing to more complex ones. Our adaptive weighting mechanism automatically manages this curriculum. It dynamically adjusts the focus between the short-term (easy) and long-term (hard) objectives based on their relative errors, allowing the model to efficiently learn a stable trajectory without the chaotic gradients from intermediate steps; however, please note that this approach can fail when excessively hard objective is adopted such as very large $M$ value in this study. The adaptive weights are computed as:
@@ 2.3 Evaluation metric
In addition to MSE, for the periodic cylinder flow problem, we evaluate the physical fidelity of the predictions using the Strouhal number ($St$). This non-dimensional quantity characterizes the dominant frequency of vortex shedding, making it a powerful metric for assessing whether the model has learned the underlying flow dynamics beyond simple numerical accuracy. To calculate it, we apply a fast Fourier transform to the $x$-velocity time series at a certain point in the cylinder’s wake. The signal is first preprocessed by removing the mean and applying a Hamming window to reduce spectral leakage. The dominant frequency, $f_{d}$, is identified from the peak of the resulting power spectrum. The Strouhal number is then computed as:
$St=\frac{f_{d}D}{U_{\infty}}$ (13)
where $D$ is the characteristic cylinder diameter and $U_{\infty}$ is the freestream velocity.
@@ 3.2 Model training details
For each of the 50 samples per PDE type, a separate standard multi-layer perceptron (MLP) model is trained, which is designed to be independent of spatial coordinates. This simple, pointwise model was deliberately chosen to test our model-agnostic temporal framework using a basic architecture with no inherent spatial inductive bias. This ensures the observed performance gains are directly attributable to the temporal methods themselves, providing a strong baseline before extending the framework to more complex, spatially-aware GNNs in Section 4. Specifically, its input is a time-series vector representing the last 30 snapshots of the solution $u$ at a single query point in the domain, and its output is the prediction for that same point’s next state. During both training and inference, the trained MLP is applied simultaneously and independently to every point in the spatial grid, effectively processing all spatial grid points in a large batch. This approach allows the model to learn a general rule for temporal evolution that can be utilized for inference at any spatial location. The final performance metrics are averaged across the 50 independent runs from 50 cases to ensure statistical robustness. The training hyperparameters are kept consistent across all experiments to facilitate a fair comparison:
@@ 3.3 Predictive performance of time integration schemes
We first evaluate the different time integration methods using a single-step rollout ($M=1$). As shown in Table 1, all schemes that predict temporal derivatives consistently outperform the conventional direct prediction method. Note that training time based on NVIDIA 3090 GPU is also shown—the same GPU machine is utilized throughout this study. And also since this study adopted lightweight AI models, inference time is estimated to be negligible, and therefore its time is not reported separately.
@@
PDE Time Integration Scheme Test MSE $\times 10^{-6}$] Training Time [s]
@@ 3.4 Application of proposed adaptive multi-step rollout
Single-step Multi-Step Rollout (MM=4)
@@ 4.1.1 Dataset and preprocessing
The dataset features a mesh scenario with a cylinder placed in a fluid flow, generating vortex shedding characterized by oscillatory flow patterns. The training mesh consists of 1,946 nodes, 11,208 edges, and 3,658 volume cells, as shown in Figure 4: this training dataset is adapted from the work of Google DeepMind [19]. The flow conditions include a maximum inlet velocity of $1.78$ m/s with a parabolic velocity distribution and a cylinder diameter of $0.074$ m. A single vortex shedding period comprises approximately 29 snapshots (or trajectories), where the time step between each is fixed at $\Delta t$=0.01 second.
For training, we use only 50 consecutive $x$-velocity snapshots. Our investigation focuses exclusively on the $x$-velocity, as this single component provides a sufficient and challenging testbed for our primary contribution: a model-agnostic temporal prediction framework. The $x$-velocity field encapsulates the most dominant physical features of the flow, including the characteristic von Kármán vortex shedding. By demonstrating our framework’s ability to accurately forecast this dynamically rich variable, we focus on a validation of the proposed temporal methods themselves. The model input consists of a sequence of $N=20$ past snapshots, and its task is to predict the next immediate snapshot.
@@ 4.1.2 Model training details
The Graph U-Net architecture employed in this study is designed with a focus on computational efficiency and simplicity, comprising only 1,177 trainable parameters (Figure 5). The encoder part of the model consists of four graph convolutional network (GCN) layers [34], with channel dimensions decreasing from 20 to 1 in the sequence 20, 15, 10, 5, and 1, where the initial dimension corresponds to the 20-channel input graph. A pooling ratio of 0.6 is applied using the gPool layer [33] after each GCN layer in the encoder, as this ratio was found to be optimal in our previous work [27]. This pooling operation reduces the number of nodes by 60% at each step, effectively capturing multi-scale features while maintaining computational tractability.
@@
In summary, the Graph U-Net model used in this section is evaluated under intentionally harsh conditions. We use a strict chronological split for our data: the model is trained on a minimal dataset of only the first 50 snapshots and then tested on its ability to forecast a long prediction horizon of the subsequent 350 snapshots. This setup, combined with the model’s limited capacity (1,177 parameters), creates a challenging 7:1 prediction-to-training ratio to rigorously assess its time-extrapolation capabilities. These constraints simulate practical engineering scenarios where only small datasets are available and computational infrastructure for model training is limited. This challenging setup rigorously tests the model’s robustness and its ability to manage error accumulation over time, particularly for complex temporal patterns like vortex shedding.
@@ 4.2.1 Predictive performance comparison
Table 3: Comparison of MSE for different time integration schemes over 350 future snapshots.
@@
Among the various approaches, Figure 6 illustrates the fundamental instability of the second-order central difference scheme. Four consecutive snapshots after 100 rollouts reveal highly oscillatory behavior: snapshots after 100 (Figure 6(a)) and 102 (Figure 6(c)) rollouts appear similar, while those after 101 (Figure 6(b)) and 103 (Figure 6(d)) exhibit similar patterns. This alternating pattern persists beyond 100 rollouts, reflecting the scheme’s fundamental limitation where $\mathbf{u}(t+\Delta t)$ depends heavily on $\mathbf{u}(t-\Delta t)$ while bypassing the intermediate state $\mathbf{u}(t)$; since $\mathbf{u}(t+\Delta t)=\mathbf{u}(t-\Delta t)+2\Delta t\cdot\frac{\delta\mathbf{u}}{\delta t}$. These results align with established findings that central difference schemes are generally unsuitable for time integration due to inherent instability.
@@
To further illustrate performance limitations, Figure 7 presents flow field predictions after 100 rollout steps for the three outperforming schemes alongside ground truth. The ground truth shows a specific vortex shedding phase behind the cylinder. All methods show noticeable discrepancies from ground truth after only 100 rollout steps. Direct prediction provides time-averaged results where flow fields remain constant during rollout progression. The corresponding error plots highlight significant prediction deviations, especially in the wake region where complex flow dynamics are most pronounced.
@@ 4.3 Extension of time integration schemes into conventional multi-step rollout scenario
We extend the previously investigated time integration schemes by incorporating conventional multi-step rollout techniques with different values of $M$, where $M$ represents the number of future snapshots considered during training. Specifically, we examine $M=1$, 22, 44, and $8$. When $M=1$, multi-step rollout is not applied, corresponding to the vanilla AR model. We adopt the conventional weighting strategy suggested by Wu et al. [30], where loss weights are set to $w_{1}=1$ for the first future step and $w_{i}=0.1$ for subsequent steps (see Eq. 8).
@@
Table 4: MSE comparison for different time integration schemes with varying multi-step rollout lengths $M$: erroneous values are shown with a gray background. Strouhal numbers (StSt) are also provided under each MSE value, where ground truth StSt value is 0.1438.
@@
($St=0.1319$) ($St=0.1387$) ($St=0.1367$) ($St=0.1455$)
Averaged time [s] 1914 2086 2633 $31$
These results reveal two key insights: First, naively increasing $M$ imposes excessive training constraints on the lightweight model (1,177 parameters), often leading to performance degradation despite the theoretical benefit of learning longer-term dependencies. Second, the time integration scheme proves more critical than the choice of MM—while most methods fail even with optimal $M$ values, Adams-Euler maintains robust performance across all $M$ values, demonstrating inherent stability for long-term AR prediction.
@@
To illustrate the improved performance achieved with Adams-Euler and $M=4$ multi-step rollout (best case in Table 4), Figure 8 shows flow field predictions after 200 and 300 rollout steps. At snapshot $t+200$ (Figures 8(a) and 8(b)), the prediction shows reasonable agreement with ground truth, capturing overall flow patterns and vortex shedding behavior. This represents substantial improvement over Figure 7, where accurate predictions were unattainable even at 100 rollout steps, clearly demonstrating the effectiveness of combining conventional multi-step rollout with the Adams-Euler scheme. However, at snapshot $t+300$ (Figures 8(c) and 8(d)), MSE increases from 0.091 to 0.105, showing noticeable divergence from ground truth: flow structure accuracy diminishes with different vortex shedding patterns. While the incorporation of multi-step rollout with Adams-Euler significantly enhances long-term prediction performance, room for improvement remains, particularly for prediction horizons extending beyond $t+200$. This motivates the development of adaptive multi-step rollout strategies explored in subsequent sections.